Betty DeRose, Inc. operates two departments, the handling department and
the packaging department. During April, the handling department reported
the following information:

                                           % complete      % complete
                                units         DM           conversion 
work in process, April 1        18,000        38%             71%
units started during April      80,000
work in process, April 30       44,000        28%             74%

The cost of beginning work in process and the costs added during April
were as follows:

                                 DM         Conversion       Total cost
work in process, April 1      $ 64,996       $152,785         $217,781
costs added during April       190,336        232,407          422,743
total costs                    255,332        385,192          640,524

Calculate the total cost of the 18,000 units in beginning work in process
using the FIFO process costing method.
Calculate the cost of goods that were started and completed in the month
of April using the FIFO process costing method.
Calculate the cost of goods completed and transferred out during April
using the FIFO process costing method.
Calculate the total cost of the handling department's work in process
inventory at April 30 using the FIFO process costing method.
Answer #1

Units started and completed = 80000-44000 = 36000 units

Equivalent units = Beginning WIP x (100% - Completion) + Units started and completed + Ending WIP x Completion
Materials = 18000 x (100%-38%) + 36000 + 44000 x 28% = 59480 units
Conversion = 18000 x (100%-71%) + 36000 + 44000 x 74% = 73780 units

Equivalent cost per unit = Cost added./ Equivalent units
Material = $190336/59480 = $3.20 per unit
Conversion = $232407/73780 = $3.15 per unit

Total cost of the 18,000 units in beginning work in process = $100708+169228 = $269936
Material = $64996 + 18000 x (100%-38%) x $3.20 = $100708
Conversion = $152785 + 18000 x (100%-71%) x $3.15 = $169228
Cost of goods that were started and completed in the month = 36000 x ($3.20+3.15) = $228600
Cost of goods completed and transferred out during April = $269936+228600 = $498536
Total cost of the handling department's work in process inventory at April 30 = $39424+102564 = $141988
Material = 44000 x 28% x $3.20 = $39424
Conversion = 44000 x 74% x $3.15 = $102564